# Image-Based Roadmaps for Vision-Only Planning and Control of Robotic Manipulators

### v2

@@ III-A Model-based Motion Planning
where $R(x)\subseteq\mathbb{R}^{3}$ is the set of all points occupied by the robot at state $x$, and $\mathcal{O}\subset\mathcal{W}$ denotes the set of obstacles within the workspace [^1] This equation only encodes collisions; joint or kinematic constraints can be incorporated in a similar manner.. Let $\mathcal{X}_{\text{free}}=\mathcal{X}\setminus\mathcal{X}_{\text{obs}}$ denote the collision-free state space. Also, let $x_{\text{start}}\in\mathcal{X}_{\text{free}}$ and $\mathcal{X}_{G}\subseteq\mathcal{X}_{\text{free}}$ represent the start state and goal region, respectively. Motion Planning Problem: Given the motion planning tuple $(\mathcal{X},R,\mathcal{W},\mathcal{U},f,x_{\text{start}},\mathcal{X}_{G})$ find a time $\mathcal{T}$ and a set of controls $u:[0,\mathcal{T}]\rightarrow\mathcal{U}$ such that the motion described by Eq. 1 satisfies $x(0)=x_{\text{start}}$, $x(\mathcal{T})\in\mathcal{X}_{G}$ and $x(t)\in\mathcal{X}_{\text{free}}$.
@@
$\footnotesize I_{\text{obs}}=\{\mathbb{K}\in I|RI(\mathbb{K})\bigcap\mathcal{IO}\neq\emptyset\}$ (3)
@@
Given the above definitions, let us define the vision-only motion planning problem.
Vision-Only Motion Planning Problem: Given the tuple $(I,RI,\mathcal{U},\mathcal{IO},Im,\mathbb{K}_{\text{start}},I_{g})$ find a time $\mathcal{T}$ and a set of controls $u:[0,\mathcal{T}]\rightarrow\mathcal{U}$ such that the motion described by $g$ satisfies $\mathbb{K}(0)=\mathbb{K}_{\text{start}}$, $\mathbb{K}(\mathcal{T})\in I_{g}$ and $\mathbb{K}(t)\in I_{\text{free}}$.
@@ IV Methodology
sample: This function typically samples the configuration space uniformly. However, in the vision-only setting, we can’t directly sample in $I$-space as we don’t have the model of the robot. Since $dim(I)>dim(\mathcal{X})$ the keypoint vectors (image state $\mathbb{K}$) that correspond to actual configurations of the robot, will lie in a lower-dimensional (equal to $dim(\mathcal{X})$) manifold in $I$-space. Thus randomly sampling $I$-space will have 00 probability of sampling a valid configuration that lies on the valid manifold [23]. We describe how to address this issues in Section IV-A, by collecting and storing valid samples directly from the real robot.
k-nearest: This function usually relies on a distance defined in $\mathcal{X}$ and finds the nearest configurations that can be connected. However, since our representation in $I$-space is now a non-holonomic system, defining this distance is very challenging, as a straight line path in $I$-space defined by a simple Euclidean distance, might not be accurately followable by a controller. To mitigate this, we describe a learning-based approach to estimate the unknown joint-distance, in Section IV-B.
coll_free: This function checks if there is a collision for an edge in $\mathcal{X}$. Again this typically requires the model $R(x)$ for the robot. In the vision only case, we propose simple yet successful method to do collision checking with an $RI(x)$ directly in $I$-space Section IV-C.
@@ IV-B1 Dataset generation for network model
While collecting the dataset of image states $\mathbb{K}$ in Section IV-A, we recorded the velocity (Eq. 5) applied to transition between consecutive frames. Using this data, we created a new dataset that includes pairs of consecutive image states ($\mathbb{K}_{1}$, $\mathbb{K}_{2}$) and their estimated joint displacements. This is calculated by multiplying the recorded velocity with the frame rate at which $\mathbb{K}$ was captured, as described in Eq. 6.
@@ V-A1 Generating Samples
We generated the required image state samples by collecting a finite dataset created by actuating the planar joints (Joints $2,4,6$) of the Franka arm, using velocities computed from Eq. 5 as explained in Section IV-A, covering the robot’s planar workspace.
@@ V-A2 Building the Roadmap
The collected samples $\mathbb{K}$ were used as nodes to build roadmaps based on the distance metrics described in Section IV-B. We refer to the roadmap using image space distances as the Image Space roadmap, and the one using learned distances as the Learned roadmap. For benchmarking, we also construct a Joint Space roadmap using actual joint displacements from encoders, used only for comparison, while preserving the model-free design. The k-neighbor value for all approaches was set to $25$. The first row of Table I shows the time required to generate each roadmap. The learned roadmap takes significantly longer to construct due to the additional computational overhead of running neural network inference during nearest neighbor computation.
@@ V-A3 Obstacle Representation and Path Planning
In the experimental setup, obstacles were modeled as virtual shapes such as a rectangle or a triangle or a circle with a safety margin, as shown in the first image of Fig. 4. Paths were generated offline for various start ($\mathbb{K}_{\text{start}}$) and goal ($\mathbb{K}_{\text{goal}}$) incorporating the obstacle avoidance logic from Section IV-C. These paths were later used in adaptive visual servoing experiments.
@@ V-B3 Planned paths for Control Experiments
We generated planned paths for two scenarios: $16$ start and goal pairs without checking for collision and $10$ pairs with collision avoidance. Paths were computed using the three roadmaps: joint space, learned, and image space. These precomputed paths were used in the control experiments described in Section V-C. The second row of Table I reports the average time each roadmap took to generate paths for these scenarios.
@@ V-C Control Experiments
The adaptive visual servoing experiments [^2] Planning and control experiments including apf videos, are available at this link. The details of how to use the link is in the supplementary ReadMe file used precomputed paths from Section V-B3, with control gains optimized to minimize rise and settling times while keeping overshoot within 5%, by careful tuning. We also performed adaptive visual-servoing experiments for the $4$ scenarios for learned and image space roadmaps as defined in Section V-A5. To highlight the advantages of planning-based control over reactive methods, we implemented an Artificial Potential Field (APF) controller [38] and tested it on the same $4$ multi-obstacle scenarios from Section V-A5.
@@
Table VII presents control performance in a proof-of-concept scenario involving simulated real-world object, as described in Section V-A6. Here, the path planned by the image space roadmap led to a collision with the object, despite a defined safety margin. Further experiments are needed to conclusively determine why the robot collided despite a collision-free planned path.
@@
In summary, both roadmaps exhibit unique benifits. The image space roadmap enables faster execution when successful but lacks reliability, while the learned roadmap offers greater robustness by leveraging joint displacement-like distances. The choice depends on the application context.
@@ VI Conclusion and Future Work
In conclusion, this work introduced a novel framework for collision-free motion planning of robotic manipulators that relied solely on visual features, eliminating the need for explicit robot models or encoder feedback. The learned roadmap offered smoother, more reliable transitions, and due to its joint displacement-based distance definition, the paths it generated maintained joint-space holonomy when it existed. In contrast, the paths produced by the image space roadmap sometimes failed to maintain joint-space holonomy, even when holonomy existed in the image space. However, the image space roadmap provided faster transient responses and simplicity, making it advantageous for applications where speed and computational efficiency were prioritized.
Future work will explore extending this approach to out-of-plane motion, incorporating real-world obstacles, and handling occlusions to develop a fully integrated control and manipulation pipeline.

### v3

@@ Abstract
Index Terms:
Motion and Path Planning, Collision Avoidance, Integrated Planning and Control
@@ III-A Model-based Motion Planning
where $R(x)\subseteq\mathbb{R}^{3}$ is the set of all points occupied by the robot at state $x$, and $\mathcal{O}\subset\mathcal{W}$ denotes the set of obstacles within the workspace [^1] This equation only encodes collisions; joint or kinematic constraints can be incorporated in a similar manner.. Let $\mathcal{X}_{\text{free}}=\mathcal{X}\setminus\mathcal{X}_{\text{obs}}$ denote the collision-free state space. Also, let $x_{\text{start}}\in\mathcal{X}_{\text{free}}$ and $\mathcal{X}_{G}\subseteq\mathcal{X}_{\text{free}}$ represent the start state and goal region, respectively.
Motion Planning Problem: Given the motion planning tuple $(\mathcal{X},R,\mathcal{W},\mathcal{U},f,x_{\text{start}},\mathcal{X}_{G})$ find a time $\mathcal{T}$ and a set of controls $u:[0,\mathcal{T}]\rightarrow\mathcal{U}$ such that the motion described by Eq. 1 satisfies $x(0)=x_{\text{start}}$, $x(\mathcal{T})\in\mathcal{X}_{G}$ and $x(t)\in\mathcal{X}_{\text{free}}$.
@@
$\footnotesize I_{\text{obs}}=\{\mathbb{K}\in I|RI(\mathbb{K})\bigcap\mathcal{IO}\neq\emptyset\}\vspace{-0.5em}$ (3)
@@
Given the above definitions, let us define the vision-only motion planning problem. Vision-Only Motion Planning Problem: Given the tuple $(I,RI,\mathcal{U},\mathcal{IO},Im,\mathbb{K}_{\text{start}},I_{g})$ find a time $\mathcal{T}$ and a set of controls $u:[0,\mathcal{T}]\rightarrow\mathcal{U}$ such that the motion described by $g$ satisfies $\mathbb{K}(0)=\mathbb{K}_{\text{start}}$, $\mathbb{K}(\mathcal{T})\in I_{g}$ and $\mathbb{K}(t)\in I_{\text{free}}$.
@@ IV Methodology
sample: This function typically samples the configuration space uniformly. However, in the vision-only setting, we can’t directly sample in $I$-space as we don’t have the model of the robot. Since $dim(I)>dim(\mathcal{X})$ the keypoint vectors (image state $\mathbb{K}$) that correspond to actual configurations of the robot, will lie in a lower-dimensional (equal to $dim(\mathcal{X})$) manifold in $I$-space. Thus randomly sampling $I$-space will have 00 probability of sampling a valid configuration that lies on the valid manifold [23]. We describe how to address this issues in Sec. IV-A, by collecting and storing valid samples directly from the real robot.
k-nearest: This function usually relies on a distance defined in $\mathcal{X}$ and finds the nearest configurations that can be connected. However, since our representation in $I$-space is now a non-holonomic system, defining this distance is very challenging, as a straight line path in $I$-space defined by a simple Euclidean distance, might not be accurately followable by a controller. To mitigate this, we describe a learning-based approach to estimate the unknown joint-distance, in Sec. IV-B.
coll_free: This function checks if there is a collision for an edge in $\mathcal{X}$. Again this typically requires the model $R(x)$ for the robot. In the vision only case, we propose simple yet successful method to do collision checking with an $RI(x)$ directly in $I$-space Sec. IV-C.
@@ IV-B1 Dataset generation for network model
While collecting the dataset of image states $\mathbb{K}$ in Sec. IV-A, we recorded the velocity (Eq. 5) applied to transition between consecutive frames. Using this data, we created a new dataset that includes pairs of consecutive image states ($\mathbb{K}_{1}$, $\mathbb{K}_{2}$) and their estimated joint displacements. This is calculated by multiplying the recorded velocity with the frame rate at which $\mathbb{K}$ was captured, as described in Eq. 6.
@@ V-A1 Generating Samples
We generated the required image state samples by collecting a finite dataset created by actuating the planar joints (Joints $2,4,6$) of the Franka arm, using velocities computed from Eq. 5 as explained in Sec. IV-A, covering the robot’s planar workspace.
@@ V-A2 Building the Roadmap
The collected samples $\mathbb{K}$ were used as nodes to build roadmaps based on the distance metrics described in Sec. IV-B. We refer to the roadmap using image space distances as the Image Space roadmap, and the one using learned distances as the Learned roadmap. For benchmarking, we also construct a Joint Space roadmap using actual joint displacements from encoders, used only for comparison, while preserving the model-free design. The k-neighbor value for all approaches was set to $25$. The first row of Table I shows the time required to generate each roadmap. The learned roadmap takes significantly longer to construct due to the additional computational overhead of running neural network inference during nearest neighbor computation.
@@ V-A3 Obstacle Representation and Path Planning
In the experimental setup, obstacles were modeled as virtual shapes such as a rectangle or a triangle or a circle with a safety margin, as shown in the first image of Fig. 4. Paths were generated offline for various start ($\mathbb{K}_{\text{start}}$) and goal ($\mathbb{K}_{\text{goal}}$) incorporating the obstacle avoidance logic from Sec. IV-C. These paths were later used in adaptive visual servoing experiments.
@@ V-B3 Planned paths for Control Experiments
We generated planned paths for two scenarios: $16$ start and goal pairs without checking for collision and $10$ pairs with collision avoidance. Paths were computed using the three roadmaps: joint space, learned, and image space. These precomputed paths were used in the control experiments described in Sec. V-C. The second row of Table I reports the average time each roadmap took to generate paths for these scenarios.
@@ V-C Control Experiments
The adaptive visual servoing experiments [^2] Planning and control experiments including apf videos, are available at this link. The details of how to use the link is in the supplementary ReadMe file. The supplementary video and relevant codes can be found in respective links used precomputed paths from Sec. V-B3, with control gains optimized to minimize rise and settling times while keeping overshoot within 5%, by careful tuning. We also performed adaptive visual-servoing experiments for the $4$ scenarios for learned and image space roadmaps as defined in Sec. V-A5. To highlight the advantages of planning-based control over reactive methods, we implemented an Artificial Potential Field (APF) controller [38] and tested it on the same $4$ multi-obstacle scenarios from Sec. V-A5.
@@
Table VII presents control performance in a proof-of-concept scenario involving simulated real-world object, as described in Sec. V-A6. Here, the path planned by the image space roadmap led to a collision with the object, despite a defined safety margin. Further experiments are needed to conclusively determine why the robot collided despite a collision-free planned path.
@@
In summary, both roadmaps exhibit unique benefits. The image space roadmap enables faster execution when successful but lacks reliability, while the learned roadmap offers greater robustness by leveraging joint displacement-like distances. The choice depends on the application context.
@@ VI Conclusion and Future Work
In conclusion, this work introduced a novel framework for collision-free motion planning of robotic manipulators that relied solely on visual features, eliminating the need for explicit robot models or encoder feedback.
The learned roadmap offered smoother, more reliable transitions, and due to its joint displacement-based distance definition, the paths it generated maintained joint-space holonomy when it existed. In contrast, the paths produced by the image space roadmap sometimes failed to maintain joint-space holonomy, even when holonomy existed in the image space. However, the image space roadmap provided faster transient responses and simplicity, making it advantageous for applications where speed and computational efficiency were prioritized.
As a future direction, to extend our approach to higher DoF and out-of-plane motion, we plan to reuse the same data collection pipeline, existing data collection pipeline while capturing trajectories that span the full 3D workspace. To ensure safe interaction with real-world obstacles, we plan to estimate monocular depth replacing the 2D safety polygon with a 3D bounding box. To improve keypoint robustness under environmental and self-occlusion, we aim to integrate a graph-based model to enforce spatial consistency within the keypoint detection network. We are also exploring the extension of this pipeline to soft and continuum robotic systems, specifically a soft origami arm. Additionally, We are also investigating a quantitative formulation to characterize and analyze the role of non-holonomic constraints in image-based motion planning.